# DecoHD: Decomposed Hyperdimensional Classification under Extreme Memory Budgets

### v1

@@ I Introduction
Hyperdimensional computing (HDC), rooted in vector-symbolic architectures (VSA), has gained traction as a lightweight and noise-tolerant paradigm for classification and learning on constrained platforms [kanerva2009hyperdimensional, imani2019framework, hernandez2021onlinehd, yun2024neurohash, yun2025hyperdimensional, yun2024hypersense, yun2025missionhd, wang2023disthd]. By encoding data into dense high-dimensional hypervectors and manipulating them with simple algebraic operations, HDC classifiers combine attractive traits: inherent robustness to device-level non-idealities, parallelism amenable to hardware acceleration, and compact compute/memory footprints. These characteristics align naturally with the demands of energy-constrained or near-memory execution environments, where classical deep learning often proves too heavy [ahmed2024internet, custode2024fast, shi2024introduction, laskaridis2024future, masum2025parahdc].
@@
A conventional HDC classifier stores one prototype hypervector per class and predicts by measuring similarity between a query and these prototypes, typically via dot product [hernandez2021onlinehd]. While this layout is simple and hardware-friendly, its storage requirement grows as $\mathcal{O}(CD)$ for $C$ classes and dimensionality $D$, which quickly dominates memory in multi-class problems. Prior work has largely reduced this footprint by compressing along the feature axis: lowering $D$, imposing sparsity, or quantizing encodings [imani2019sparsehd, morris2019comphd, ge2020classification, verges2025classification, imani2019quanthd]. Although effective in reducing memory and multiply–accumulate cost, such strategies erode the concentration-of-measure properties that stabilize similarity in high dimensions. At extreme budgets, they degrade both robustness to analog noise [liang2024robust, liang2024stridehd] and classification accuracy.
@@
This tension is especially acute in deployment scenarios where memory is the first-order constraint. Across emerging AI-on-edge applications—such as TinyML-based sensing and recognition [celen2025tinymltraffic, li2025tinymlpositioning, liu2025solarml], batteryless IoT platforms [custode2024fast, tabrizchi2025intermittent], and federated/distributed AI in heterogeneous systems [dudziak2022fedoras, zhu2024shufflefl, chu2025priprune]—models must be aggressively compressed while retaining resilience to device-level imperfections. In these regimes, preserving $D$ while reducing the number of stored dense hypervectors offers a more principled pathway than shrinking dimensionality.
@@
Inspired by decomposition techniques successful in compact deep neural networks—such as low-rank adaptation, modular fine-tuning, and shared-codebook representations [yin2021towards, yin2022batude, horvath2023maestro, tastan2025loft, kwon2023tinytrain]—we propose a complementary direction tailored to HDC. Rather than compressing hypervector dimensionality, we reduce the number of stored dense hypervectors by parameterizing class prototypes as structured compositions of a small, shared set of channels, combined using HDC-native binding ($\otimes$) and weighted bundling ($\oplus$), as illustrated in Figure 1. Concretely, $N$ layers expose $\{L_{i}\}$ learnable channels; selecting one channel per layer yields $M=\prod_{i}L_{i}$ bound paths. A lightweight class-specific head $W\in\mathbb{R}^{C\times M}$ aggregates these paths to form logits. This design scales memory roughly with $\sum_{i}L_{i}$, while maintaining full-$D$ representations and their robustness benefits. This decomposition also introduces a tunable memory–compute trade-off: increasing the number of layers $N$ or channels reduces storage but raises test-time computation, since more bound-path compositions must be evaluated; deeper factorizations with small $L$ provide larger representational capacity at the cost of higher inference cost, a favorable exchange in regimes where memory is the primary constraint.
@@
Naively decomposing dense prototypes post hoc is unstable in high dimensions, since binding is multiplicative and orderless, making the inverse factorization ill-posed [frady2020resonator, poduval2024hdqmf]. Instead, we introduce DecoHD, which learns directly in the decomposed space. Low-dimensional latents are expanded into channel hypervectors via fixed random projections, and training optimizes the bind–bundle pipeline end-to-end using cross-entropy. At inference, the method relies solely on $\otimes$, $\oplus$, and dot products, ensuring compatibility with existing HDC/VSA accelerators and memory-centric hardware paths.
@@
In summary, DecoHD tackles the memory bottleneck of conventional HDC by decomposing class prototypes into a compact set of shared channels while preserving the ambient dimensionality $D$. This design retains the stability and robustness of high-dimensional representations under extreme budgets and aligns with the needs of edge-AI deployments across TinyML, federated learning, and memory-centric accelerators, where compact yet reliable models are critical [laskaridis2024future]. Our evaluation shows that DecoHD achieves accuracy comparable to conventional HDC while delivering up to $277\times$ higher energy efficiency and $35\times$ faster inference than CPU baselines, $14\times$ and $3.7\times$ improvements over GPU, and $2.0\times$ and $2.4\times$ gains over a conventional HDC ASIC, all with only $0.38\times$ memory.
@@
  3. 3.
    Efficiency and deployment. We analyze memory and compute complexity of the stacked binding design ($M{=}\prod_{i}L_{i}$ bound paths from $\sum_{i}L_{i}$ channels) and show that DecoHD supports a pure-HDC inference flow using only $\otimes$, $\oplus$, and dot products. An ASIC implementation achieves up to $277\times$ energy efficiency and $35\times$ speedup at $0.38\times$ memory usage compared to the baseline, underscoring suitability for near-memory accelerators, TinyML, and federated/edge systems under tight memory and energy budgets.
@@ II-A HDC and Vector Symbolic Architectures
Hyperdimensional computing—also known as vector symbolic architectures—represents symbols and compositional structure with high-dimensional hypervectors manipulated by binding ($\otimes$), bundling ($\oplus$), and permutation ($\pi$[kanerva2009hyperdimensional]. Classification typically superposes encoded examples into per-class prototypes and scores by dot product. This operation set maps well to emerging in/near-memory platforms and low-power edge contexts: recent work demonstrates robust HDC pipelines under voltage scaling and binary encodings [liang2024stridehd], specialized VSA/HDC designs for efficiency [moghadam2025idvsa], and accelerator paths that preserve the native HDC interface [masum2025parahdc]. At the system level, the push toward consumer edge-AI [laskaridis2024future] and intermittently powered/batteryless devices [ahmed2024internet, custode2024fast, tabrizchi2025intermittent] further motivates lightweight, memory-frugal models. Concurrently, the embedded-systems community highlights in/near-memory directions [shi2024introduction] and 3D integration trends [yin2021towards, yin2022batude] that favor linear, data-parallel HDC kernels. Our design intentionally keeps HDC’s native bind–bundle–dot scoring to remain drop-in compatible with such accelerators.
@@ II-B Model-Size Reduction for HDC-based Classifiers
Many efficiency efforts in HDC implicitly shrink along the feature axis—e.g., binary/low-precision encodings or simplified encoders—which reduce stored bits and arithmetic but also reduce effective dimensionality, weakening high-$D$ averaging that stabilizes similarity and increasing sensitivity to perturbations under tight power/voltage budgets [liang2024stridehd, zhang2021assessing]. In contrast, DecoHD preserves the ambient $D$ and compresses orthogonally along the class axis by sharing a small pool of per-layer channels across classes and learning only the light bundling head. This reduces the # of dense hypervectors without altering the bind–bundle–score pipeline, and complements systems work on efficient HDC datapaths and accelerators [masum2025parahdc, moghadam2025idvsa].
@@ II-C Decomposition in DNNs and HDC for Model Compression
Decomposition is a standard route to shrink DNNs while preserving input/output dimensionality—e.g., trainable low-rank/tensor factorizations and related adapters [horvath2023maestro, tastan2025loft]. Our approach brings this spirit to HDC but avoids ill-posed post hoc factorization of dense class prototypes. Instead, DecoHD learns in a decomposed parameterization: a small, shared set of per-layer channels whose stacked bindings yield $M$ path features, followed by class-wise weighted bundling. This keeps $D$ unchanged, retains native HDC operations ($\otimes,\oplus$), and shifts storage from $C\times D$ prototypes to compact channels plus a light bundling head—well-suited to near-memory execution and resource-constrained edge deployment [shi2024introduction, laskaridis2024future, ahmed2024internet]. Finally, our end-to-end training of low-dimensional latents (expanded by fixed random projections) aligns with recent trends in efficient/edge training and personalization [kwon2023tinytrain, dudziak2022fedoras, zhu2024shufflefl].
@@ IV Experiments
    ISOLET [isolet_54] 617 26 6,238 1,559 Voice recognition
    UCIHAR [human_activity_recognition_using_smartphones_240] 261 12 6,213 1,554 Activity recognition (mobile)
    PAMAP2 [pamap2_physical_activity_monitoring_231] 75 5 611,142 101,582 Activity recognition (IMU)
    PAGE [page_blocks_classification_78] 10 5 4,925 548 Page layout blocks classification
@@ IV-A Experimental Setup
    Target memory budgets. Results are reported under budgets $\leq m\in(0,1]$ and denoted in parentheses as $(\leq m)$, where $m$ is the model size relative to a conventional HDC table with $C\times D$ parameters (at $b$ bits each). DecoHD replaces this table with a lightweight bundling head $W\!\in\!\mathbb{R}^{C\times M}$ and a bank of channel latents totaling $L_{\mathrm{tot}}=\sum_{i}L_{i}$ of size $d$, yielding the approximate normalized footprint
    $m\;\approx\;\frac{CM+L_{\mathrm{tot}}D}{CD},$
    with $M=\prod_{i}L_{i}$. Selecting $(N,\{L_{i}\},D)$ to satisfy $m$ enforces the target budget.
@@ V Conclusions
    We presented DecoHD, a class-axis decomposition for HDC that yields memory-lean inference. Across benchmarks, it matches a strong non-reduced HDC baseline closely (average accuracy gap $\approx$0.1–0.15%), improves robustness to random bit-flip noise, and reduces trainable parameters by up to $\sim$97% at saturation. In hardware, an ASIC realization delivers up to $277\times$ higher energy efficiency at $0.38\times$ memory with consistent speedups over CPU/GPU and a conventional HDC ASIC. Depth reduces memory but increases test-time computation and can reduce accuracy under fixed bound paths; our findings suggest selecting the shallowest factorization that meets the memory budget to preserve throughput and accuracy.
@@ Acknowledgements
    This work was supported in part by the DARPA Young Faculty Award, the National Science Foundation (NSF) under Grants #2431561, #2127780, #2319198, #2321840, #2312517, and #2235472, the Semiconductor Research Corporation (SRC), the Office of Naval Research through the Young Investigator Program Award, and Grants #N00014-21-1-2225 and #N00014-22-1-2067, and Army Research Office Grant #W911NF2410360. Additionally, support was provided by the Air Force Office of Scientific Research under Award #FA9550-22-1-0253, along with generous gifts from Xilinx and Cisco.

### v2

@@ I Introduction
Hyperdimensional computing (HDC), rooted in vector-symbolic architectures (VSA), has gained traction as a lightweight and noise-tolerant paradigm for classification and learning on constrained platforms [14, 12, 9, 38, 37, 35, 36, 32]. By encoding data into dense high-dimensional hypervectors and manipulating them with simple algebraic operations, HDC classifiers combine attractive traits: inherent robustness to device-level non-idealities, parallelism amenable to hardware acceleration, and compact compute/memory footprints. These characteristics align naturally with the demands of energy-constrained or near-memory execution environments, where classical deep learning often proves too heavy [1, 5, 28, 16, 22].
@@
A conventional HDC classifier stores one prototype hypervector per class and predicts by measuring similarity between a query and these prototypes, typically via dot product [9]. While this layout is simple and hardware-friendly, its storage requirement grows as $\mathcal{O}(CD)$ for $C$ classes and dimensionality $D$, which quickly dominates memory in multi-class problems. Prior work has largely reduced this footprint by compressing along the feature axis: lowering $D$, imposing sparsity, or quantizing encodings [13, 24, 8, 31, 11]. Although effective in reducing memory and multiply–accumulate cost, such strategies erode the concentration-of-measure properties that stabilize similarity in high dimensions. At extreme budgets, they degrade both robustness to analog noise [18, 19] and classification accuracy.
@@
This tension is especially acute in deployment scenarios where memory is the first-order constraint. Across emerging AI-on-edge applications—such as TinyML-based sensing and recognition [2, 17, 20], batteryless IoT platforms [5, 29], and federated/distributed AI in heterogeneous systems [6, 40, 3]—models must be aggressively compressed while retaining resilience to device-level imperfections. In these regimes, preserving $D$ maintains the concentration-of-measure properties central to HDC, making class-axis compression a more principled pathway than shrinking dimensionality.
@@
Inspired by decomposition techniques successful in compact deep neural networks—such as low-rank adaptation, modular fine-tuning, and shared-codebook representations [34, 33, 10, 30, 15]—we propose a complementary direction tailored to HDC. Rather than compressing hypervector dimensionality, we reduce the number of stored dense hypervectors by parameterizing class prototypes as structured compositions of a small, shared set of channels, combined using HDC-native binding ($\otimes$) and weighted bundling ($\oplus$), as illustrated in Figure 1. Concretely, $N$ layers expose $\{L_{i}\}$ learnable channels; selecting one channel per layer yields $M=\prod_{i}L_{i}$ bound paths. A lightweight class-specific head $W\in\mathbb{R}^{C\times M}$ aggregates these paths to form logits. This design scales memory roughly with $\sum_{i}L_{i}$, while maintaining full-$D$ representations and their robustness benefits. This decomposition also introduces a tunable memory–compute trade-off: increasing the number of layers $N$ or channels reduces storage but raises test-time computation, since more bound-path compositions must be evaluated; deeper factorizations with small $L$ provide larger representational capacity at the cost of higher inference cost, a favorable exchange in regimes where memory is the primary constraint.
@@
Naively decomposing dense prototypes post hoc is unstable in high dimensions, since binding is multiplicative and orderless, making the inverse factorization ill-posed [7, 25]. Instead, we introduce DecoHD, which learns directly in the decomposed space. Low-dimensional latents are expanded into channel hypervectors via fixed random projections, and training optimizes the bind–bundle pipeline end-to-end using cross-entropy. At inference, the method relies solely on $\otimes$, $\oplus$, and dot products, ensuring compatibility with existing HDC/VSA accelerators and memory-centric hardware paths.
@@
In summary, DecoHD tackles the memory bottleneck of conventional HDC by decomposing class prototypes into a compact set of shared channels while preserving the ambient dimensionality $D$. This design retains the stability and robustness of high-dimensional representations under extreme budgets and aligns with the needs of edge-AI deployments across TinyML, federated learning, and memory-centric accelerators, where compact yet reliable models are critical [16]. Our evaluation shows that DecoHD achieves accuracy comparable to conventional HDC while delivering up to $277\times$ higher energy efficiency and $35\times$ faster inference than CPU baselines, $14\times$ and $3.7\times$ improvements over GPU, and $2.0\times$ and $2.4\times$ gains over a conventional HDC ASIC, all with only $0.38\times$ memory.
@@ II-A HDC and Vector Symbolic Architectures
Hyperdimensional computing—also known as vector symbolic architectures—represents symbols and compositional structure with high-dimensional hypervectors manipulated by binding ($\otimes$), bundling ($\oplus$), and permutation ($\pi$[14]. Classification typically superposes encoded examples into per-class prototypes and scores by dot product. This operation set maps well to emerging in/near-memory platforms and low-power edge contexts: recent work demonstrates robust HDC pipelines under voltage scaling and binary encodings [19], specialized VSA/HDC designs for efficiency [23], and accelerator paths that preserve the native HDC interface [22]. At the system level, the push toward consumer edge-AI [16] and intermittently powered/batteryless devices [1, 5, 29] further motivates lightweight, memory-frugal models. Concurrently, the embedded-systems community highlights in/near-memory directions [28] and 3D integration trends [34, 33] that favor linear, data-parallel HDC kernels. Our design intentionally keeps HDC’s native bind–bundle–dot scoring to remain drop-in compatible with such accelerators.
@@ II-B Model-Size Reduction for HDC-based Classifiers
Many efficiency efforts in HDC implicitly shrink along the feature axis—e.g., binary/low-precision encodings or simplified encoders—which reduce stored bits and arithmetic but also reduce effective dimensionality, weakening high-$D$ averaging that stabilizes similarity and increasing sensitivity to perturbations under tight power/voltage budgets [19, 39]. In contrast, DecoHD preserves the ambient $D$ and compresses orthogonally along the class axis by sharing a small pool of per-layer channels across classes and learning only the light bundling head. This reduces the # of dense hypervectors without altering the bind–bundle–score pipeline, and complements systems work on efficient HDC datapaths and accelerators [22, 23].
@@ II-C Decomposition in DNNs and HDC for Model Compression
Decomposition is a standard route to shrink DNNs while preserving input/output dimensionality—e.g., trainable low-rank/tensor factorizations and related adapters [10, 30]. Our approach brings this spirit to HDC but avoids ill-posed post hoc factorization of dense class prototypes. Instead, DecoHD learns in a decomposed parameterization: a small, shared set of per-layer channels whose stacked bindings yield $M$ path features, followed by class-wise weighted bundling. This keeps $D$ unchanged, retains native HDC operations ($\otimes,\oplus$), and shifts storage from $C\times D$ prototypes to compact channels plus a light bundling head—well-suited to near-memory execution and resource-constrained edge deployment [28, 16, 1]. Finally, our end-to-end training of low-dimensional latents (expanded by fixed random projections) aligns with recent trends in efficient/edge training and personalization [15, 6, 40].
@@ IV Experiments
    ISOLET [4] 617 26 6,238 1,559 Voice recognition
    UCIHAR [27] 261 12 6,213 1,554 Activity recognition (mobile)
    PAMAP2 [26] 75 5 611,142 101,582 Activity recognition (IMU)
    PAGE [21] 10 5 4,925 548 Page layout blocks classification
@@ IV-A Experimental Setup
    Target memory budgets. Results are reported under budgets $\leq m\in(0,1]$ and denoted in parentheses as $(\leq m)$, where $m$ is the model size relative to a conventional HDC table with $C\times D$ parameters (at $b$ bits each). DecoHD replaces this table with a lightweight bundling head $W\!\in\!\mathbb{R}^{C\times M}$ and a bank of channel latents totaling $L_{\mathrm{tot}}=\sum_{i}L_{i}$ of size $d$, yielding the approximate normalized footprint $m\;\approx\;\frac{CM+L_{\mathrm{tot}}D}{CD},$ with $M=\prod_{i}L_{i}$. Selecting $(N,\{L_{i}\},D)$ to satisfy $m$ enforces the target budget.
@@ V Conclusions
    We presented DecoHD, a gradient-based class-axis decomposition for HDC that enables memory-lean inference under extreme resource constraints. Across benchmarks, it closely matches a strong non-reduced HDC baseline (average accuracy gap $\approx$0.1–0.15%), improves robustness to random bit-flip noise, and reduces trainable parameters by up to $\sim$97% at saturation. In hardware, an ASIC realization achieves up to $277\times$ higher energy efficiency at $0.38\times$ memory, with consistent speedups over CPU/GPU and a conventional HDC ASIC. By preserving full hypervector dimensionality while aggressively reducing stored prototypes, DecoHD is well suited for TinyML and edge deployments in memory-first and near-memory systems. Depth reduces memory but increases test-time computation; thus, the shallowest factorization that meets the target memory budget is preferred to preserve throughput and accuracy.
@@ Acknowledgements
    This work was supported in part by the DARPA Young Faculty Award, the National Science Foundation (NSF) under Grants #2431561, #2127780, #2319198, #2321840, #2312517, and #2235472, the Semiconductor Research Corporation (SRC), the Office of Naval Research through the Young Investigator Program Award and Grants #N00014-21-1-2225 and #N00014-22-1-2067, Army Research Office Grant #W911NF2410360, the U.S. Army Combat Capabilities Development Command Army Research Laboratory under Support Agreement No. USMA 21050, and DARPA under Support Agreement No. USMA 23004. Additionally, support was provided by the Air Force Office of Scientific Research under Award #FA9550-22-1-0253, along with generous gifts from Xilinx and Cisco.